# mode = titleGraph Filtering for Multimodal Recommendation

### v2

@@ 4.2.1 Graph Construction for User–Item Interactions
Similarly as in existing GF-based recommendation methods Park et al. (2024, 2025); Kim et al. (2025), the item–item similarity graph for user–item interactions is constructed as
@@
where $\alpha$ is a hyperparameter controlling the asymmetric normalization along users and items Park et al. (2024, 2025); Kim et al. (2025). Additionally, we adjust $\tilde{P}$ using the Hadamard power $\bar{P}=\tilde{P}^{\circ s}$ as properly adjusting the item–item similarity graph using the Hadamard power was shown to produce more accurate recommendations Park et al. (2024, 2025); Kim et al. (2025), where $\circ$ denotes the Hadamard (element-wise) power and $s$ is a filter adjustment hyperparameter.
@@ 4.2.2 Graph Construction for Textual and Visual Modalities
If we naïvely apply the same graph construction strategy used for user–item interactions to the textual and visual modalities, then we may fail due to the existence of outliers or negative values in feature vectors generated through sentence-transformers Reimers and Gurevych (2019) and pre-trained convolutional neural networks (CNNs) He and McAuley (2016a). As illustrated in Figure 4, constructing item–item similarity graphs for multimodal information using this strategy can lead to two critical issues. First, outliers in the feature vectors may cause certain elements of the graph to take excessively large values, thereby dominating the similarity graph and suppressing the contribution of other similarities, which are then forced to remain close to zero (see Figure 4(a)). This distortion makes it difficult to accurately capture the true relationships between items. Second, when feature vectors include both positive and negative values, the row and column sums in the normalization process can approach zero (see Figure 4(b)). Dividing by such near-zero values results in singularities in the normalized graph, ultimately leading to instability and unreliable graph structures. Therefore, it is necessary to construct a graph structure in the form of similarity graphs tailored for multimodal information. While there exist a variety of similarity calculation strategies such as the Pearson correlation coefficient Cohen et al. (2009) and the Gaussian kernel Shuman et al. (2013), we empirically found that cosine similarity-based calculation guarantees superior performance on benchmark datasets in terms of recommendation accuracy consistently. In this context, we adopt cosine similarity to construct item–item similarity graphs unless otherwise stated.
To be specific, following Zhang et al. (2021); Zhou and Shen (2023); Yu et al. (2023), MM-GF constructs an item–item similarity graph by applying kkNN sparsification on raw features corresponding to each modality $m$. Given $N$ items, the similarity score $S_{i,j}^{m}$ between item $i$ and $j$ is computed using cosine similarity based on their raw modality-$m$ feature vectors $(x_{i}^{m},x_{j}^{m})$, as follows:
@@
Here, $S_{i,j}^{m}$ represents the element located at the $i$-th row and $j$-th column of the similarity matrix $S^{m}\in\mathbb{R}^{N\times N}$. To sparsify the graph, we apply $k$-nearest neighbor preservation in Zhou and Shen (2023) to convert the weighted matrix $S^{m}$ into an unweighted version. For each item $i$, only its top-$k$ most similar neighbors are kept, resulting in the binary matrix $\hat{S}^{m}$ defined by:
@@
where $D_{m,\text{r}}=\operatorname{diag}(\hat{S}^{m}\mathbf{1})$ and $D_{m,\text{c}}=\operatorname{diag}(\mathbf{1}^{T}\hat{S}^{m})$ are the diagonal degree matrices for modality $m$. Likewise, we adjust $\tilde{P}_{m}$ as $\bar{P}_{m}=\tilde{P}_{m}^{\circ s}$ in the graph construction process. [^1] While using different $\alpha$’s and $s$’s for each modality certainly increases recommendation accuracy, we use the same values of $\alpha$ and $s$ as those for user–item interactions across multiple modalities for simplicity.
@@ 4.2.3 Filter Design with Multiple Modalities
Next, we are interested in how to precisely adjust the frequency bound of item–item similarity graphs. Specifically, when we adopt polynomial graph filters in Park et al. (2024, 2025); Kim et al. (2025), it is necessary that all eigenvalues of the item–item similarity graph lie within $[0,1]$. This approach defines a polynomial graph filter as $\sum_{k=1}^{K}a_{k}\bar{P}_{\star}^{k}$, which corresponds to the frequency response function $h(\lambda)=\sum_{k=1}^{K}a_{k}(1-\lambda)^{k}$. However, as we observed in Figure  2 of Section 4.1, certain eigenvalues exceed the specified bounds.
@@
While directly computing the coefficients $a_{k}$ in Eq. (11) provides precise control over the frequency response, it also restricts flexibility because the coefficients are fixed in advance. To enhance flexibility, we treat $a_{k}$ as tunable hyperparameters, allowing data-driven selection through validation rather than adhering to a fixed analytic form.
@@ 5 Experimental Results and Analyses
Table 2: The statistics of three benchmark datasets.
Dataset # of users # of items # of interactions Sparsity
Baby 19,445 7,050 160,792 99.88%
Sports 35,598 18,357 296,337 99.95%
Clothing 39,387 23,033 278,677 99.97%
@@
  • RQ6 : How sensitive is MM-GF to variations in its hyperparameters?
@@ 5.1 Experimental Settings
Datasets. We use three widely used benchmark datasets in recent MRSs Zhang et al. (2021); Yu et al. (2023); Zhou and Shen (2023); Zhou et al. (2023a), which were collected from Amazon McAuley et al. (2015): (a) Baby, (b) Sports and Outdoors, and (c) Clothing, Shoes, and Jewelry, which we refer to as Baby, Sports, and Clothing, respectively, in brief. [^3] Datasets are officially available at https://jmcauley.ucsd.edu/data/amazon/links.html. The above three datasets contain textual and visual features as well as user–item interactions. For the textual modality, we use 384-dimensional textual embeddings by combining the title, descriptions, categories, and brand of each item and adopt sentence-transformers Reimers and Gurevych (2019). For the visual modality, we use 4,096-dimensional visual embeddings obtained by applying pre-trained CNNs He and McAuley (2016a). Table 1 provides a summary of the statistics for each dataset.
@@
Competitors. To validate the effectiveness of MM-GF, we conduct a comparative analysis with seven state-of-the-art recommendation methods, especially those built upon neural network models including GCNs. The benchmark methods include not only a single-modal recommendation method (LightGCN He et al. (2020)) but also multimodal recommendation methods (VBPR He and McAuley (2016b), GRCN Wei et al. (2020), LATTICE Zhang et al. (2021), BM3 Zhou et al. (2023b), FREEDOM Zhou and Shen (2023), and MGCN Yu et al. (2023)).
Table 3: Performance comparison among MM-GF and competitors. The best and second-best performers are highlighted in bold and underline, respectively.
@@
Method Recall@10 NDCG@10 Recall@20 NDCG@20 Recall@10 NDCG@10 Recall@20 NDCG@20 Recall@10 NDCG@10 Recall@20 NDCG@20
@@
Evaluation protocols. We use the same dataset split for training, test, and validation sets, as well as the textual and visual data processing for feature extraction, as those in previous studies Zhang et al. (2021); Yu et al. (2023); Zhou and Shen (2023). To assess the top-$K$ recommendation performance, we adopt the widely used metrics from prior studies He et al. (2020); He and McAuley (2016b); Wei et al. (2020); Zhang et al. (2021); Zhou et al. (2023b); Zhou and Shen (2023); Yu et al. (2023); Shen et al. (2021); Park et al. (2024), namely recall and normalized discounted cumulative gain (NDCG), where $K\in\left\{{10,20}\right\}$. From RQ2 to RQ6 excecpt RQ4, we report the results only for NDCG@$20$ as similar trends are observed for other metrics.
Implementation details. For a fair comparison, we implement the proposed MM-GF method and all benchmark methods using MMRec Zhou (2023), an open-sourced multimodal recommendation framework. [^4] The toolbox is available at https://github.com/enoche/MMRec. Unless otherwise stated, for MM-GF, the best-performing hyperparameters $(a_{1},a_{2},a_{3},\beta,\gamma)$ in Eqs. (11) and (15) are set to (-0.7, 1.1, -0.1, 0.1, 0), (0.7, 1.8, 1.5, 1.9, 0.5), and (1.9, 0.7, -0.1, 0.5, 0.2) for the Baby, Sports, and Clothing datasets, respectively, using the validation set. Additionally, we apply the proposed polynomial graph filter to the user–item interaction matrix, restricting the expansion to the third order (i.e., $K=3$ in Eq. (11)). For the textual and visual modalities, we adopt a linear LPF, specified as $\bar{P}_{m}^{f}=\bar{P}_{m}-\lambda_{min}^{m}I$, for each modality $m$. These design choices are primarily guided by empirical evidence. The parameter $k$ in Eq. (9) was empirically set to 20 in our experiments. All experiments are carried out on a machine with Intel (R) 12-Core (TM) i7-9700K CPUs @ 3.60 GHz and an NVIDIA GeForce RTX A6000 GPU.
Table 4: Runtime comparison among MM-GF and representative GCN-based competitors. The best performer is highlighted in bold. NDCG refers to NDCG@2020.
@@
Method NDCG Time NDCG Time NDCG Time
@@
5.2 Recommendation Accuracy (RQ1)
  1. (i)
  2. (ii)
  3. (iii)
5.3 Runtime and Scalability (RQ2)
Table 4 summarizes the runtime of MM-GF and GCN-based competitors that perform well (GRCN, LATTICE, BM3, and MGCN) on the three datasets. For the GCN-based methods, runtime refers to the training time, whereas, for the GF-based method (MM-GF), it indicates the processing time, as in Shen et al. (2021); Park et al. (2024). We observe that, on the Baby dataset, MM-GF performs approximately $\times 100.4$ faster than MGCN, which is the best GCN-based multimodal recommendation method, while exhibiting even higher recommendation accuracy. This is because MM-GF operates solely on straightforward matrix calculations without a costly training process. A similar tendency is observed on other two datasets. In consequence, the proposed MM-GF method is advantageous in terms of runtime as well as recommendation accuracy.
@@ 5.3 Runtime and Scalability (RQ2)
Dataset $M$-GF $M$-GF-t $M$-GF-tv $M$-GF-pf $M$-GF-f
Baby 0.0522 0.0448 0.0448 0.0484 0.0420
Sports 0.0593 0.0493 0.0485 0.0562 0.0486
Clothing 0.0513 0.0377 0.0348 0.0445 0.0308

### v3

@@ 4.2.1 Graph Construction for User–Item Interactions
Similarly as in existing GF-based recommendation methods DBLP:conf/sigir/ParkSS24; DBLP:conf/www/ParkYS25; DBLP:conf/www/KimCPS25, the item–item similarity graph for user–item interactions is constructed as
@@
where $\alpha$ is a hyperparameter controlling the asymmetric normalization along users and items DBLP:conf/sigir/ParkSS24; DBLP:conf/www/ParkYS25; DBLP:conf/www/KimCPS25. Additionally, we adjust $\tilde{P}$ using the Hadamard power $\bar{P}=\tilde{P}^{\circ s}$ as properly adjusting the item–item similarity graph using the Hadamard power was shown to produce more accurate recommendations DBLP:conf/sigir/ParkSS24; DBLP:conf/www/ParkYS25; DBLP:conf/www/KimCPS25, where $\circ$ denotes the Hadamard (element-wise) power and $s$ is a filter adjustment hyperparameter.
@@ 4.2.2 Graph Construction for Textual and Visual Modalities
If we naïvely apply the same graph construction strategy used for user–item interactions to the textual and visual modalities, then we may fail due to the existence of outliers or negative values in feature vectors generated through sentence-transformers DBLP:conf/emnlp/ReimersG19 and pre-trained convolutional neural networks (CNNs) DBLP:conf/www/HeM16. As illustrated in Figure 4, constructing item–item similarity graphs for multimodal information using this strategy can lead to two critical issues. First, outliers in the feature vectors may cause certain elements of the graph to take excessively large values, thereby dominating the similarity graph and suppressing the contribution of other similarities, which are then forced to remain close to zero (see Figure 4(a)). This distortion makes it difficult to accurately capture the true relationships between items. Second, when feature vectors include both positive and negative values, the row and column sums in the normalization process can approach zero (see Figure 4(b)). Dividing by such near-zero values results in singularities in the normalized graph, ultimately leading to instability and unreliable graph structures. Therefore, it is necessary to construct a graph structure in the form of similarity graphs tailored for multimodal information. While there exist a variety of similarity calculation strategies such as the Pearson correlation coefficient cohen2009pearson and the Gaussian kernel DBLP:journals/spm/ShumanNFOV13, we empirically found that cosine similarity-based calculation guarantees superior performance on benchmark datasets in terms of recommendation accuracy consistently. In this context, we adopt cosine similarity to construct item–item similarity graphs unless otherwise stated.
To be specific, following DBLP:conf/mm/Zhang00WWW21; DBLP:conf/mm/ZhouS23; DBLP:conf/mm/Yu0LB23, MM-GF constructs an item–item similarity graph by applying kkNN sparsification on raw features corresponding to each modality $m$. Given $N$ items, the similarity score $S_{i,j}^{m}$ between item $i$ and $j$ is computed using cosine similarity based on their raw modality-$m$ feature vectors $(x_{i}^{m},x_{j}^{m})$, as follows:
@@
Here, $S_{i,j}^{m}$ represents the element located at the $i$-th row and $j$-th column of the similarity matrix $S^{m}\in\mathbb{R}^{N\times N}$. To sparsify the graph, we apply $k$-nearest neighbor preservation in DBLP:conf/mm/ZhouS23 to convert the weighted matrix $S^{m}$ into an unweighted version. For each item $i$, only its top-$k$ most similar neighbors are kept, resulting in the binary matrix $\hat{S}^{m}$ defined by:
@@
where $D_{m,\text{r}}=\operatorname{diag}(\hat{S}^{m}\mathbf{1})$ and $D_{m,\text{c}}=\operatorname{diag}(\mathbf{1}^{T}\hat{S}^{m})$ are the diagonal degree matrices for modality $m$. Likewise, we adjust $\tilde{P}_{m}$ as $\bar{P}_{m}=\tilde{P}_{m}^{\circ s}$ in the graph construction process. [^1] While using different $\alpha$’s and $s$’s for each modality certainly increases recommendation accuracy, we use the same values of $\alpha$ and $s$ as those for user–item interactions across multiple modalities for simplicity. Consequently, we construct item–-item similarity graphs for the textual and visual modalities, denoted by $\bar{P}_{\text{txt}}$ and $\bar{P}_{\text{img}}$, respectively.
@@ 4.2.3 Filter Design with Multiple Modalities
Next, we are interested in how to precisely adjust the frequency bound of item–item similarity graphs. Specifically, when we adopt polynomial graph filters in DBLP:conf/sigir/ParkSS24; DBLP:conf/www/ParkYS25; DBLP:conf/www/KimCPS25, it is necessary that all eigenvalues of the item–item similarity graph lie within $[0,1]$. This approach defines a polynomial graph filter as $\sum_{k=1}^{K}a_{k}\bar{P}_{\star}^{k}$, which corresponds to the frequency response function $h(\lambda)=\sum_{k=1}^{K}a_{k}(1-\lambda)^{k}$. However, as we observed in Figure  2 of Section 4.1, certain eigenvalues exceed the specified bounds.
@@
Intuitively, Eq. (11) shifts and rescales the spectrum of $\bar{P}_{\star}$ so that all eigenvalues fall into the interval $0,\lambda^{\star}]$, and then applies a polynomial graph filter on this transformed spectrum, thereby resolving the instability that occurs when conventional polynomial graph filters are directly applied to matrices whose eigenvalues lie outside the assumed bounded range.
While directly specifying the coefficients $a_{k}$ in Eq. (11) provides precise, analytically designed control over the frequency response, it also limits expressiveness because the filter shape is fixed a priori by the designer. In practice, since it is often difficult to know in advance which frequency components should be emphasized or attenuated for a given dataset, a hand-crafted filter may fail to align with the true spectral characteristics of real-world recommendation graphs. To address this, we treat $a_{k}$ as tunable hyperparameters, allowing data-driven selection via validation rather than enforcing a single fixed analytic form.
@@
4.3 Computational Complexity Analysis
Our proposed method leverages training-free graph filtering with a polynomial filter design and spectral components derived from the item–item graph. Here, we analyze the computational complexity of the proposed MM-GF method from three viewpoints.
Polynomial graph filtering.
The polynomial graph filter of order $K$ is implemented via $K$ matrix-vector multiplications. Given an $N\times N$ item-item graph (where $N$ is the number of items), each multiplication has a complexity of $\mathcal{O}(N^{2})$. Thus, the overall filtering step has a time complexity of $\mathcal{O}(KN^{2})$, which is efficient for moderate values of $K$.
Spectral component computation.
In our implementation, we use exact eigendecomposition to obtain the graph Laplacian’s spectrum, which has a computational complexity of $\mathcal{O}(N^{3})$. However, it is important to emphasize that in our experiments—conducted on standard multimodal recommendation benchmarks—the graph sizes are moderate, and the eigendecomposition step does not dominate the overall runtime, accounting for only a small fraction of the total computation. Hence, the runtime is practically acceptable, as also supported by our empirical timing results (see Table 4).
Scalability via approximation.
For large-scale applications where full eigendecomposition may become infeasible, our framework remains applicable by relying on approximate methods. Specifically, since the polynomial filter only requires the largest and smallest eigenvalues of the Laplacian, we can employ efficient approximation techniques such as the Lanczos algorithm cullum2002lanczos to estimate these extremal eigenvalues with significantly lower complexity than $\mathcal{O}(N^{3})$ arora2005fast. Thus, our method maintains scalability without compromising its training-free and model-agnostic design.
@@ 5 Experimental Results and Analyses
  • RQ6: How does the polynomial graph filter of MM-GF operate in the spectral domain?
  • RQ7: How sensitive is MM-GF to variations in its hyperparameters?
@@ 5.1 Experimental Settings
Datasets. We use three widely used benchmark datasets in recent MRSs DBLP:conf/mm/Zhang00WWW21; DBLP:conf/mm/Yu0LB23; DBLP:conf/mm/ZhouS23; zhou2023comprehensive, which were collected from Amazon DBLP:conf/sigir/McAuleyTSH15: (a) Baby, (b) Sports and Outdoors, and (c) Clothing, Shoes, and Jewelry, which we refer to as Baby, Sports, and Clothing, respectively, in brief. [^3] Datasets are officially available at https://jmcauley.ucsd.edu/data/amazon/links.html. The above three datasets contain textual and visual features as well as user–item interactions. For the textual modality, we use 384-dimensional textual embeddings by combining the title, descriptions, categories, and brand of each item and adopt sentence-transformers DBLP:conf/emnlp/ReimersG19. For the visual modality, we use 4,096-dimensional visual embeddings obtained by applying pre-trained CNNs DBLP:conf/www/HeM16. Table 2 provides a summary of the statistics for each dataset.
Table 2: The statistics of three benchmark datasets.
Dataset # of users # of items # of interactions Sparsity
Baby 19,445 7,050 160,792 99.88%
Sports 35,598 18,357 296,337 99.95%
Clothing 39,387 23,033 278,677 99.97%
@@
Competitors. To validate the effectiveness of MM-GF, we conduct a comparative analysis with seven state-of-the-art recommendation methods, especially those built upon neural network models including GCNs. The benchmark methods include not only a single-modal recommendation method (LightGCN DBLP:conf/sigir/0001DWLZ020) but also multimodal recommendation methods (VBPR DBLP:conf/aaai/HeM16, GRCN DBLP:conf/mm/WeiWN0C20, LATTICE DBLP:conf/mm/Zhang00WWW21, BM3 DBLP:conf/www/ZhouZLZMWYJ23, FREEDOM DBLP:conf/mm/ZhouS23, and MGCN DBLP:conf/mm/Yu0LB23).
Evaluation protocols. Since MM-GF does not require parameter training, the notion of a conventional training set does not apply. Nevertheless, for consistency with widely adopted evaluation protocols, we randomly partition each user’s interaction history by using 80% of interactions to construct the graph, 10% for validation, and the remaining 10% for testing. To assess the top-$K$ recommendation performance, we adopt the widely used metrics from prior studies DBLP:conf/sigir/0001DWLZ020; DBLP:conf/aaai/HeM16; DBLP:conf/mm/WeiWN0C20; DBLP:conf/mm/Zhang00WWW21; DBLP:conf/www/ZhouZLZMWYJ23; DBLP:conf/mm/ZhouS23; DBLP:conf/mm/Yu0LB23; DBLP:conf/cikm/ShenWZSZLL21; DBLP:conf/sigir/ParkSS24, namely recall and normalized discounted cumulative gain (NDCG), where $K\in\left\{{10,20}\right\}$. From RQ2 to RQ6 excecpt RQ4, we report the results only for NDCG@$20$ as similar trends are observed for other metrics. Note that MM-GF is fully deterministic; once the data split and hyperparameters are fixed, it produces identical predictions across runs. Therefore, we do not report statistics over multiple random seeds or conduct run-wise significance tests for MM-GF, and instead compare single-run results that are directly determined by the chosen configuration.
Implementation details. For a fair comparison, we implement the proposed MM-GF method and all benchmark methods using MMRec DBLP:conf/mmasia/Zhou23, an open-sourced multimodal recommendation framework. [^4] The toolbox is available at https://github.com/enoche/MMRec. To calculate frequency bound of item–item similarity graphs, the minimum and maximum eigenvalues $(\lambda_{min},\lambda_{max})$ are computed once per dataset and modality using the torch.linalg.eigvalsh routine. Unless otherwise stated, for MM-GF, the best-performing hyperparameters $(a_{1},a_{2},a_{3},\beta,\gamma)$ in Eqs. (11) and (15) are selected on the validation set from the following ranges: for the multimodal aggregation $P+\beta P_{\text{text}}+\gamma P_{\text{img}}$, we tune $\beta$ and $\gamma$ in $[0,2]$ with a step size of $0.1$; for the polynomial filter, we search $a_{1},a_{2},a_{3}$ in [2,2][-2,2] with a step size of $0.1$. The finally chosen values are (0.1,1.1,0.7,0.1,0)(-0.1,1.1,-0.7,0.1,0), $(1.5,1.8,0.7,1.9,0.5)$, and (0.1,0.7,1.9,0.5,0.2)(-0.1,0.7,1.9,0.5,0.2) for the Baby, Sports, and Clothing datasets, respectively. Additionally, we apply the proposed polynomial graph filter to the user–item interaction matrix, restricting the expansion to the third order (i.e., $K=3$ in Eq. (11)). For the textual and visual modalities, we adopt a linear LPF, specified as $\bar{P}_{m}^{f}=\bar{P}_{m}-\lambda_{min}^{m}I$, for each modality $m$. These design choices are primarily guided by empirical evidence. The parameter $k$ in Eq. (9) was empirically set to 20 in our experiments. All experiments are carried out on a machine with Intel (R) 12-Core (TM) i7-9700K CPUs @ 3.60 GHz and an NVIDIA GeForce RTX A6000 GPU.
5.2 Recommendation Accuracy (RQ1)
  1. (i)
  2. (ii)
  3. (iii)
Table 3: Performance comparison among MM-GF and competitors. The best and second-best performers are highlighted in bold and underline, respectively. Here, R@$K$ and N@$K$ indicates Recall@$K$ and NDCG@$K$, respectively. We further conduct a user-level paired Wilcoxon signed-rank test, confirming that MM-GF significantly outperforms the strongest baseline (i.e., MGCN) on all evaluated datasets and metrics ($p\ll 0.01$).
@@ 5.2 Recommendation Accuracy (RQ1)
Method R@10 N@10 R@20 N@20 R@10 N@10 R@20 N@20 R@10 N@10 R@20 N@20
@@
5.3 Runtime and Scalability (RQ2)
Table 4 summarizes the runtime of MM-GF and GCN-based competitors that perform well (GRCN, LATTICE, BM3, and MGCN) on the three datasets. For the GCN-based methods, runtime refers to the training time, whereas, for the GF-based method (MM-GF), it indicates the processing time, as in DBLP:conf/cikm/ShenWZSZLL21; DBLP:conf/sigir/ParkSS24. We observe that, on the Baby dataset, MM-GF performs approximately $\times 100.4$ faster than MGCN, which is the best GCN-based multimodal recommendation method, while exhibiting even higher recommendation accuracy. This is because MM-GF operates solely on straightforward matrix calculations without a costly training process. A similar tendency is observed on other two datasets. In consequence, the proposed MM-GF method is advantageous in terms of runtime as well as recommendation accuracy.
Table 4: Runtime comparison among MM-GF and representative GCN-based competitors. The best performer is highlighted in bold. N@20 refers to NDCG@2020.
@@ 5.3 Runtime and Scalability (RQ2)
Method N@20 Time N@20 Time N@20 Time
@@
Dataset $M$-GF $M$-GF-t $M$-GF-tv $M$-GF-pf $M$-GF-f $M$-GF-h
Baby 0.0522 0.0448 $(\downarrow 14.2\%)$ 0.0448 $(\downarrow 14.2\%)$ 0.0484 $(\downarrow 7.3\%)$ 0.0420 $(\downarrow 19.5\%)$ 0.0417 $(\downarrow 20.1\%)$
Sports 0.0593 0.0493 $(\downarrow 16.9\%)$ 0.0485 $(\downarrow 18.2\%)$ 0.0562 $(\downarrow 5.2\%)$ 0.0486 $(\downarrow 18.0\%)$ 0.0429 $(\downarrow 27.7\%)$
Clothing 0.0513 0.0377 $(\downarrow 26.5\%)$ 0.0348 $(\downarrow 32.2\%)$ 0.0445 $(\downarrow 13.3\%)$ 0.0308 $(\downarrow 40.0\%)$ 0.0394 $(\downarrow 23.2\%)$